# Finite-Time Analysis of Simultaneous Double Q-learning

### v1

@@ 1 Introduction
Reinforcement learning (RL) is a class of learning algorithms for finding an optimal policy in unknown environments through experiences [1]. $Q$-learning [2] is one of the most popular RL algorithms, which finds an optimal policy by estimating the optimal $Q$-function. Its effectiveness has been established through many successful applications [3, 4, 5, 6] as well as theoretical convergence analysis [7, 8, 9, 10, 11, 12, 13].
@@
Despite its empirical successes and theoretical achievements, $Q$-learning is known to suffer from overestimation in the $Q$-estimator, known as the maximization bias [1], which slows down the convergence of the corresponding policy to an optimal policy especially in environments with a large number of actions and heterogeneous action spaces. To overcome this obstacle, the so-called double $Q$-learning was proposed in [14], which empirically proved that the maximization bias can be reduced by using double $Q$-estimators instead of the single $Q$-estimator. Since its introduction, double $Q$-learning has been successfully applied in practice [15, 16, 17], and analyzed thoroughly in [18, 19]. However, from a practical standpoint, double $Q$-learning introduces an additional random selection step between the two $Q$-estimators, which can slow down its convergence speed compared to standard $Q$-learning despite its advantage in mitigating maximization bias.
Motivated by the aforementioned discussion, this paper proposes a modified double $Q$-learning called simultaneous double $Q$-learning (SDQ), which departs from the original in two key aspects: 1) Elimination of random selection: It dispenses with the need for random selection between the two $Q$-estimators, a step that can slow down convergence in the original double $Q$-learning. Through experiments, effectiveness of SDQ is demonstrated in this paper; 2) Different roles of $Q$-estimators: The role of each $Q$-estimator is distinct from those employed in the traditional double $Q$-learning algorithm in the sense that the roles of the two $Q$-estimators are swapped in the update rule. This change allows us to view SDQ through the lens of switching system models [11, 12, 13], enabling more efficient finite-time analysis. Moreover, we provide expected finite-time error bounds of SDQ instead of error bounds with high probability constraints [18, 19] in the literature. To our knowledge, this is the first such analysis for double $Q$-learning.
@@
$\text{max}\{\mathbb{E}[\lVert Q_{k}^{A}-Q^{*}\rVert_{\infty}],\,\mathbb{E}[\lVert Q_{k}^{B}-Q^{*}\rVert_{\infty}]\}$ $\leq\frac{120\alpha^{1/2}\lvert S\times A\rvert}{d_{\text{min}}^{9/2}(1-\gamma)^{11/2}}+\frac{48\rho^{k-4}k^{4}\lvert S\times A\rvert^{3/2}}{(1-\gamma)},$
where $\lvert S\times A\rvert$ is the number of the state-action pairs, $\gamma$ is the discount factor, $d_{\text{min}}$ is the minimum state-action occupation frequency, $\alpha\in(0,1)$ is the constant step-size, $Q^{*}$ is a vectorized optimal $Q$-function, $Q^{A}_{k}$ and $Q^{B}_{k}$ are its double estimates at the current time $k$, and $\rho\coloneqq 1-\alpha d_{\text{min}}(1-\gamma)\in(0,1)$ is the exponential decay rate.
Although the switching system model has been first introduced in [11, 12, 13], we extend this view to double $Q$-learning and provide a new finite-time analysis. We note that this extension is not trivial due to more complex structures of double $Q$-learning, which includes an extra error term between the two $Q$-estimators that makes the analysis much more challenging compared to the standard $Q$-learning. Therefore, the techniques used in the previous studies cannot be directly applied to double $Q$-learning. In this paper, new approaches have been developed to overcome this challenge. Details on the proposed analysis can be found in Section 5, 6. Finally, the main contributions are summarized as follows:
@@
    Simultaneous double $Q$-learning (SDQ) is proposed to address maximization bias while exhibiting favorable convergence properties. Moreover, this modification enables viewing double $Q$-learning through the lens of a switching system, enabling more efficient finite-time analysis;
@@ 2 Related Works
There have been relatively little researches on the convergence analysis of double $Q$-learning. The first convergence proof of double $Q$-learning has been provided by [14], but it only established asymptotic convergence. Finite-time convergence results have been recently given in [18, 19]. Both conducted convergence analysis on synchronous and asynchronous double $Q$-learning with non-i.i.d. (independent and identically distributed) observation model. In this paper, we analyze the convergence properties of a variant of double $Q$-learning where the $Q$-estimators serve different roles compared to the original approach. We specifically focus on asynchronous double $Q$-learning with modifications and i.i.d. observation model for our analysis. Regarding the step-size, the previous analysis in [18, 19] imposes more restrictive ranges of step-sizes for their analysis. However, the proposed analysis allows more flexible step-sizes $\alpha\in(0,1)$. Moreover, the existing studies have derived error bounds with high probability constraints. However, the proposed results are presented in terms of the expected error bounds, which provide additional insights and complement the existing analysis.
@@ 3.1 Markov Decision Problem
We consider the infinite-horizon discounted Markov decision problem (MDP), where the agent sequentially takes actions to maximize cumulative discounted rewards. In a Markov decision process with the state-space ${\cal S}:=\{1,2,\ldots,|{\cal S}|\}$ and action-space ${\cal A}:=\{1,2,\ldots,|{\cal A}|\}$, the decision maker selects an action $a\in{\cal A}$ with the current state $s$, then the state transits to a state $s^{\prime}$ with probability $P(s,a,s^{\prime})$, and the transition incurs a reward $r(s,a,s^{\prime})$, where $P(s,a,s^{\prime})$ is the state transition probability from the current state $s\in{\cal S}$ to the next state $s^{\prime}\in{\cal S}$ under action $a\in{\cal A}$, and $r(s,a,s^{\prime})$ is the reward function. For convenience, we consider a deterministic reward function and simply write $r(s_{k},a_{k},s_{k+1})=:r_{k},k\in\{0,1,\ldots\}$.
A deterministic policy, $\pi:{\cal S}\to{\cal A}$, maps a state $s\in{\cal S}$ to an action $\pi(s)\in{\cal A}$. The objective of the MDP is to find a deterministic optimal policy, $\pi^{*}$, such that the cumulative discounted rewards over infinite time horizons is maximized, i.e.,
$\pi^{*}:=\operatorname*{arg\,max}_{\pi\in\Theta}{\mathbb{E}}\left[\left.\sum_{k=0}^{\infty}{\gamma^{k}r_{k}}\right|\pi\right],$
where $\gamma\in[0,1)$ is the discount factor, $\Theta$ is the set of all admissible deterministic policies, $(s_{0},a_{0},s_{1},a_{1},\ldots)$ is a state-action trajectory generated by the Markov chain under policy $\pi$, and ${\mathbb{E}}[\cdot|\pi]$ is an expectation conditioned on the policy $\pi$. The $Q$-function under policy $\pi$ is defined as
$Q^{\pi}(s,a)={\mathbb{E}}\left[\left.\sum_{k=0}^{\infty}{\gamma^{k}r_{k}}\right|s_{0}=s,a_{0}=a,\pi\right],\quad s\in{\cal S},a\in{\cal A},$
and the optimal $Q$-function is defined as $Q^{*}(s,a)=Q^{\pi^{*}}(s,a)$ for all $s\in{\cal S},a\in{\cal A}$. Once $Q^{*}$ is known, then an optimal policy can be retrieved by the greedy policy $\pi^{*}(s)=\operatorname*{arg\,max}_{a\in{\cal A}}Q^{*}(s,a)$. Throughout, we assume that the MDP is ergodic so that the stationary state distribution exists and the MDP is well posed.
@@ 3.2 Switching system
Since the switching system [20, 21] is a special form of nonlinear systems [22], we first consider the nonlinear system
@@
where $x_{k}\in{\mathbb{R}}^{n}$ is the state and $f:{\mathbb{R}}^{n}\to{\mathbb{R}}^{n}$ is a nonlinear mapping. An important concept in dealing with the nonlinear system is the equilibrium point. A point $x=x^{*}$ in the state-space is said to be an equilibrium point of (1) if it has the property that whenever the state of the system starts at $x^{*}$, it will remain at $x^{*}$ [22]. For (1), the equilibrium points are the real roots of the equation $f(x)=x$. The equilibrium point $x^{*}$ is said to be globally asymptotically stable if for any initial state $x_{0}\in{\mathbb{R}}^{n}$, $x_{k}\to x^{*}$ as $k\to\infty$.
Next, let us consider the particular system, called the linear switching system,
@@
where $x_{k}\in{\mathbb{R}}^{n}$ is the state, $\sigma\in{\mathcal{M}}:=\{1,2,\ldots,M\}$ is called the mode, $\sigma_{k}\in{\mathcal{M}}$ is called the switching signal, and $\{A_{\sigma},\sigma\in{\mathcal{M}}\}$ are called the subsystem matrices. The switching signal can be either arbitrary or controlled by the user under a certain switching policy. Especially, a state-feedback switching policy is denoted by $\sigma_{k}=\sigma(x_{k})$. A more general class of systems is the affine switching system
@@
where $b_{\sigma_{k}}\in{\mathbb{R}}^{n}$ is the additional input vector, which also switches according to $\sigma_{k}$. Due to the additional input $b_{\sigma_{k}}$, its stabilization becomes much more challenging.
@@ 3.3 Double Q-learning
Double $Q$-learning [14] is a variant of $Q$-learning [2], which can overcome the maximization bias in its update by updating one of the two $Q$-estimators $Q^{A}_{k}$ and $Q^{B}_{k}$, which is selected randomly. Therefore, the corresponding update can be presented as follows:
$Q_{k+1}^{A}(s_{k},a_{k})$ $=\zeta_{k}Q_{k}^{A}(s_{k},a_{k})+\alpha_{k}\zeta_{k}\{r_{k}+\gamma Q_{k}^{B}(s_{k+1},\text{argmax}_{a\in A}Q_{k}^{A}(s_{k},a))-Q_{k}^{A}(s_{k},a_{k})\},$
$Q_{k+1}^{B}(s_{k},a_{k})$ $=(1-\zeta_{k})Q_{k}^{B}(s_{k},a_{k})+\alpha_{k}(1-\zeta_{k})\{r_{k}+\gamma Q_{k}^{A}(s_{k+1},\text{argmax}_{a\in B}Q_{k}^{B}(s_{k},a))$
$\quad-Q_{k}^{B}(s_{k},a_{k})\},$ (3)
where the updates for $Q^{A}_{k}$ and $Q^{B}_{k}$ are chosen randomly with the Bernoulli random variable $\zeta_{k}$ which satisfies $\mathbb{P}(\zeta_{k}=0)=\mathbb{P}(\zeta_{k}=1)=0.5$. By eliminating the max operator in its updates, it is known to reduce effectively the maximization bias.

### v2

@@ Abstract
keywords:
Simultaneous double QQ‐learning , double QQ‐learning , finite‐time analysis , maximization bias , switching systems
journal: Neurocomputing
@@ 1 Introduction
Reinforcement learning (RL) is a class of learning algorithms for finding an optimal policy in unknown environments through interactions with the environment [1]. Among them, $Q$-learning [2] is one of the most widely studied and practically successful methods, which aims to learn an optimal policy by iteratively estimating the optimal action–value function. Owing to its simplicity and model-free nature, $Q$-learning has been successfully applied to a wide range of problems, including control, robotics, and game playing [3, 4, 5, 6]. From a theoretical perspective, its convergence properties have also been extensively studied under various setting, such as stochastic approximation frameworks and finite-time anlayses [7, 8, 9, 10, 11, 12, 13].
@@
Despite its empirical successes and theoretical achievements, $Q$-learning is known to suffer from overestimation in the $Q$-estimator, known as the maximization bias [1]. This bias arises because the $Q$-value update selects the maximum action-value estimate, often leading to overestimation due to noise in the sampled estimates. For instance, when multiple actions are available, even small overestimations can accumulate through repeated updates, systematically skewing the $Q$-function. This issue becomes particularly severe in environments with a large number of actions or heterogeneous action spaces, where it can significantly slow the convergence of the policy to an optimal solution. To overcome this obstacle, the so-called double $Q$-learning was proposed in [14], which empirically demonstrated that the maximization bias can be reduced by using double $Q$-estimators instead of the single $Q$-estimator. Since its introduction, double $Q$-learning has been successfully applied in practice [15, 16, 17], and analyzed thoroughly in [18, 19]. However, from a practical standpoint, double $Q$-learning employs a random switching mechanism between two $Q$-estimators to mitigate maximization bias. While this mechanism effectively reduces overestimation, it relies on an alternating update scheme between the two $Q$-estimators. As a result, the overall learning process can theoretically take up to twice as long to converge under the same step-size settings [30].
Motivated by the aforementioned discussion, this paper proposes a modified double $Q$-learning called simultaneous double $Q$-learning (SDQ), which departs from the original in two key aspects: 1) Elimination of random selection: It dispenses with the need for random selection between the two $Q$-estimators, a step that can slow down convergence in the original double $Q$-learning. This design replaces the stochastic estimator-selection mechanism in double $Q$-learning with a simultaneous update scheme, where both estimators are updated concurrently at each iteration. This eliminates the randomness in estimator selection, and the update structure becomes deterministic, which aligns naturally with the switching-system interpretation adopted in our theoretical framework. 2) Different roles of $Q$-estimators: In the original double $Q$-learning, the two estimators play asymmetric roles: one estimator selects the greedy action based on its own values, while the other provides the target for the update. In contrast, SDQ introduces a cross-referenced mechanism in which each estimator uses the other to determine the greedy action, but computes the target value using its own estimate. Specifically, the updates of SDQ can be expressed as:
$Q_{k+1}^{A}(s_{k},a_{k})$ $=Q_{k}^{A}(s_{k},a_{k})+\alpha_{k}\{r_{k+1}+\gamma Q_{k}^{A}(s_{k+1},\text{argmax}_{a\in\mathcal{A}}Q_{k}^{B}(s_{k+1},a))-Q_{k}^{A}(s_{k},a_{k})\},$
$Q_{k+1}^{B}(s_{k},a_{k})$ $=Q_{k}^{B}(s_{k},a_{k})+\alpha_{k}\{r_{k+1}+\gamma Q_{k}^{B}(s_{k+1},\text{argmax}_{a\in\mathcal{A}}Q_{k}^{A}(s_{k+1},a))-Q_{k}^{B}(s_{k},a_{k})\},$
where $Q_{k}^{A}$ and $Q_{k}^{B}$ denote two separate estimators of the optimal action–value function $Q^{*}$ at iteration $k$. The pair $(s_{k},a_{k})\in\mathcal{S}\times\mathcal{A}$ represents the state–action pair sampled at time $k$, $r_{k+1}$ is the immediate reward observed after taking action $a_{k}$ at state $s_{k}$, and $s_{k+1}$ is the subsequent state. The scalar $\alpha_{k}>0$ denotes the step size, and $\gamma\in(0,1)$ is the discount factor. Each estimator updates itself using the greedy action determined by the other estimator, while evaluating the target value with its own estimate. This mutual role exchange creates a symmetric interaction between the two estimators, and the resulting update equations form a coupled pair that can be naturally modeled as a discrete-time switching system [20]. Such a symmetric formulation provides an analytical structure that facilitates finite-time convergence analysis.
@@
$\max\bigl\{\mathbb{E}\lVert Q_{k}^{A}-Q^{*}\rVert_{\infty},\;\mathbb{E}\lVert Q_{k}^{B}-Q^{*}\rVert_{\infty}\bigr\}$ $\leq\frac{120\,\alpha^{1/2}\lvert\mathcal{S}\times\mathcal{A}\rvert}{d_{\min}^{9/2}(1-\gamma)^{11/2}}+\,\frac{48\,\rho^{\,k-4}k^{4}\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{1-\gamma}\,$, (1)
where $\lvert\mathcal{S}\times\mathcal{A}\rvert$ is the number of the state-action pairs, $d_{\text{min}}$ is the minimum state-action occupation frequency, $\alpha\in(0,1)$ is the constant step-size and $\rho\coloneqq 1-\alpha d_{\text{min}}(1-\gamma)\in(0,1)$ is the exponential decay rate.
Although the switching system model has been first introduced in [11, 12, 13], we extend this view to double $Q$-learning and provide a new finite-time analysis. We note that this extension is not trivial because the two estimators are coupled through their update rules. These additional dependencies complicate the finite-time analysis compared to standard $Q$-learning. Therefore, the techniques used in the previous studies cannot be directly applied to double $Q$-learning. In this paper, new approaches have been developed to overcome this challenge. Details on the proposed analysis can be found in Section 5, 6. Finally, the main contributions are summarized as follows:
@@
    SDQ is proposed to address maximization bias while exhibiting favorable convergence properties. Moreover, this modification enables double $Q$-learning to be viewed through the lens of a switching system and enables more efficient finite-time analysis.
@@ 2 Related works
There has been a growing body of research on finite-time analyses of $Q$-learning and its variants. Beyond double $Q$-learning, several recent studies have investigated finite-sample guarantees for vanilla $Q$-learning. For instance, [24] developed a Lyapunov-based theory for Markovian stochastic approximation. This theory provides finite-sample bounds for asynchronous tabular $Q$-learning under Markovian sampling.
Other works [25, 26] have also established non-asymptotic error bounds for $Q$-learning with similar Markovian sampling assumptions.
There have been relatively few studies on the convergence analysis of double $Q$-learning. The first convergence proof was provided by [14], but it only established asymptotic convergence. Finite-time convergence results have been more recently presented in [18, 19], which analyzed both synchronous and asynchronous double $Q$-learning under non-i.i.d. observation models. Unlike prior analyses of $Q$-learning and double $Q$-learning that typically rely on Markovian sampling and cover-time conditions, our SDQ adopts an i.i.d. stochastic exploration assumption in which each state-action pair pair is independently drawn from a stationary distribution that is positive for all state-action pair. This formulation simplifies the finite-time analysis while preserving the essential stochastic nature of RL. We note that the Markovian setting is more practical and realistic, as it captures temporal correlations commonly observed in RL environments. However, with modest additional effort, such as incorporating mixing-time or cover-time conditions, the proposed framework can also be extended to the Markovian case as demonstrated in [34]. Regarding the step-size, the previous analyses in [18, 19] impose more restrictive ranges for convergence, whereas the proposed analysis allows a broader class of step-sizes $\alpha\in(0,1)$.
Moreover, existing works in [18, 19] have primarily established high-probability error bounds, which provide probabilistic guarantees on the learning process. In contrast, our analysis focuses on expected error bounds, which characterize the expected estimation accuracy and offer a complementary viewpoint. While the two types of results serve different purposes, they are closely related: an expected error bound can typically be converted into a probabilistic one through concentration inequalities. Therefore, the expected bound used in this study should not be viewed as weaker or stronger, but rather as a complementary formulation that provides mean-error characterization aligned with our system-theoretic analysis framework. In this sense, our result complements the existing probabilistic analyses and contributes to a more complete understanding of finite-time behavior in double $Q$-learning.
@@ 3.1 Markov decision problem
We focus on an infinite-horizon discounted Markov decision process (MDP) in which an agent learns an optimal policy by maximizing the expected discounted sum of future rewards through sequential interactions with the environment. The environment is modeled by a finite state space ${\cal S}:=\{1,2,\ldots,|{\cal S}|\}$ and a finite action space ${\cal A}:=\{1,2,\ldots,|{\cal A}|\}$, where $|{\cal S}|$ and $|{\cal A}|$ denote the cardinalities of the state and action spaces, respectively. At each step, given the current state $s\in{\cal S}$, the agent chooses an action $a\in{\cal A}$, and the system transitions to a next state $s^{\prime}\in{\cal S}$ with probability $P(s^{\prime}\mid s,a)$. It receives a reward $r(s,a,s^{\prime})$. For simplicity, we assume that the reward function is deterministic and denote it by $r(s_{k},a_{k},s_{k+1})=:r_{k+1}$, where $k\in\{0,1,\ldots\}$. A deterministic policy $\pi:\mathcal{S}\to\mathcal{A}$ assigns to each state $s\in\mathcal{S}$ a specific action $\pi(s)\in\mathcal{A}$. The objective of the Markov decision problem is to determine an optimal policy $\pi^{*}$ that maximizes the expected cumulative discounted rewards over an infinite horizon:
$\pi^{*}:=\arg\max_{\pi\in\Theta}\mathbb{E}\!\left[\sum_{k=0}^{\infty}\gamma^{k}r_{k+1}\,\bigg|\,\pi\right],$
where $\gamma\in[0,1)$ is the discount factor, $\Theta$ denotes the set of all admissible deterministic policies, $(s_{0},a_{0},s_{1},a_{1},\ldots)$ represents a state–action trajectory generated under policy $\pi$, and $\mathbb{E}[\cdot|\pi]$ indicates the expectation conditioned on $\pi$. The $Q$-function associated with a policy $\pi$ is defined as
$Q^{\pi}(s,a)=\mathbb{E}\!\left[\sum_{k=0}^{\infty}\gamma^{k}r_{k+1}\,\bigg|\,s_{0}=s,\,a_{0}=a,\,\pi\right],\quad(s,a)\in\mathcal{S}\times\mathcal{A}.$
The optimal $Q$-function is given by $Q^{*}(s,a)=Q^{\pi^{*}}(s,a)$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$. Once $Q^{*}$ is obtained, the optimal policy can be recovered via the greedy rule:
$\pi^{*}(s)=\arg\max_{a\in\mathcal{A}}Q^{*}(s,a).$
Throughout the paper, we assume that the MDP is ergodic. It ensures the existence of a stationary state distribution and the well-posedness of the problem.
@@ 3.2 Switching system
Following standard notions in control theory [20, 21, 22], a discrete-time switching system can be regarded as a particular instance of a nonlinear dynamical system. We briefly revisit this concept, as it forms the analytical foundation for representing the update mechanism of $Q$-learning. We begin with a general nonlinear discrete-time system:
@@
where $x_{k}\in\mathbb{R}^{n}$ denotes the system state and $f:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is a nonlinear mapping. A point $x^{*}\in\mathbb{R}^{n}$ is called an equilibrium point of (2) if the state remains at $x^{*}$ whenever the system starts from $x_{0}=x^{*}$. For (2), equilibrium points are the real roots of the equation $f(x)=x$. Moreover, an equilibrium $x^{*}$ is said to be globally asymptotically stable if, for any initial condition $x_{0}\in\mathbb{R}^{n}$, the state trajectory satisfies $x_{k}\to x^{*}$ as $k\to\infty$.
A subclass of nonlinear systems is the linear switching system [20], expressed as
@@
where $x_{k}\in\mathbb{R}^{n}$ is the state, $\sigma_{k}\in\mathcal{M}:=\{1,2,\ldots,M\}$ denotes the mode at time $k$, and $\{A_{\sigma}\}_{\sigma\in\mathcal{M}}$ are the subsystem matrices. The switching signal $\sigma_{k}$ may vary arbitrarily or follow a prescribed policy, such as a state-feedback rule $\sigma_{k}=\sigma(x_{k})$. A more general formulation is the affine switching system:
@@
where $b_{\sigma_{k}}\in\mathbb{R}^{n}$ represents a mode-dependent additional input vector. The presence of this additional affine term generally increases the difficulty of ensuring system stability.
@@ 3.3 Double Q-learning
Double $Q$-learning [14] is a variant of $Q$-learning [2], which can reduce the maximization bias in its update by updating one of the two $Q$-estimators $Q^{A}_{k}$ and $Q^{B}_{k}$, which is selected randomly. Therefore, the corresponding update can be presented as follows:
$Q_{k+1}^{A}(s_{k},a_{k})$ $=\zeta_{k}Q_{k}^{A}(s_{k},a_{k})+\alpha_{k}\zeta_{k}\{r_{k+1}+\gamma Q_{k}^{B}(s_{k+1},\text{argmax}_{a\in\mathcal{A}}Q_{k}^{A}(s_{k+1},a))-Q_{k}^{A}(s_{k},a_{k})\},$
$Q_{k+1}^{B}(s_{k},a_{k})$ $=(1-\zeta_{k})Q_{k}^{B}(s_{k},a_{k})+\alpha_{k}(1-\zeta_{k})\{r_{k+1}+\gamma Q_{k}^{A}(s_{k+1},\text{argmax}_{a\in\mathcal{A}}Q_{k}^{B}(s_{k+1},a))-Q_{k}^{B}(s_{k},a_{k})\},$ (4)
where $Q_{k}^{A}$ and $Q_{k}^{B}$ denote two separate estimators of the optimal action–value function $Q^{*}$ at iteration $k$. The pair $(s_{k},a_{k})\in\mathcal{S}\times\mathcal{A}$ represents the state–action pair sampled at time $k$, $r_{k+1}$ is the immediate reward observed after taking action $a_{k}$ at state $s_{k}$, and $s_{k+1}$ is the subsequent state. The scalar $\alpha_{k}>0$ denotes the step size at iteration $k$, and $\gamma\in(0,1)$ is the discount factor. The Bernoulli random variable $\zeta_{k}\in\{0,1\}$ determines which estimator is updated at iteration $k$, with $\mathbb{P}(\zeta_{k}=0)=\mathbb{P}(\zeta_{k}=1)=0.5$. At each iteration, only one of the two estimators is updated using the greedy action determined by the other estimator. By eliminating the max operator in its updates, it is known to reduce effectively the maximization bias.